The student travels by car for 60 minutes. In the first 18 minutes, the car crossed a certain part of the road at an average road speed of 11 m / s. Find the average road speed during the remaining 42 minutes of travel, assuming that the average road speed along the entire journey was 21 m / s.


1
Expert's answer
2021-03-04T11:56:53-0500
"v_{av}=\frac{v_1t_1+v_2t_2}{t_1+t_2}\\21=\frac{(11)(18)+42v_2}{60}\\v_2=25.3\frac{m}{s}"